Each month, the amount of simple interest earned in an account that varies directly with the starting amount in the account. Suppose that a starting amount of $7100 earns $426 interest in a month. what is the starting amount that earns $510 interest in a month?
Answer by rfer(16322)

7100/426=x/510
426x=3621000
x=3621000/426
x=$8500